OpenStudy (nophunintended):

determine which ordered pair is a solution of y = -5x+10 (-15,5) (2,0) (-1,-5) (3,25) on which of the following lines does the point (7,1) lie? y-5x = 4 y = x-10 y= =4x+3 y = -x+8 Ill fan and medal, thanks

OpenStudy (anonymous):

like a first step y= -5x +10 factoriz out the -5 so what will get ?

OpenStudy (anonymous):

will get y = -5(x-2) yes ?

OpenStudy (anonymous):

so do you can solce it for x ?

OpenStudy (anonymous):

solve -sorry

OpenStudy (anonymous):

please collaborate

OpenStudy (nophunintended):

I'm sorry for the late response and such but I don't fully understand how to solve for x @jhonyy9

OpenStudy (anonymous):

just make the y=0 what mean that will get -5x +10 = 0 factorized out the -5 how i have made it above aou get -5(x+2) = 0 so when will be these equation equal zero ?

OpenStudy (anonymous):

you get - sorry

OpenStudy (nophunintended):

I found the answer but thanks for the help @jhonyy9

OpenStudy (anonymous):

np was my pleasure good luck bye bye